### v1

# Architectural Co-Design for Zero-Shot Anomaly Detection: Decoupling Representation and Dynamically Fusing Features in CLIP

Abstract
Pre-trained Vision-Language Models (VLMs) face a significant adaptation gap when applied to Zero-Shot Anomaly Detection (ZSAD), stemming from their lack of local inductive biases for dense prediction and their reliance on inflexible feature fusion paradigms. We address these limitations through an Architectural Co-Design framework that jointly refines feature representation and cross-modal fusion. Our method integrates a parameter-efficient Convolutional Low-Rank Adaptation (Conv-LoRA) adapter to inject local inductive biases for fine-grained representation, and introduces a Dynamic Fusion Gateway (DFG) that leverages visual context to adaptively modulate text prompts, enabling a powerful bidirectional fusion. Extensive experiments on diverse industrial and medical benchmarks demonstrate superior accuracy and robustness, validating that this synergistic co-design is critical for robustly adapting foundation models to dense perception tasks.
Index Terms—  anomaly detection, multimodal feature fusion, vision-language model, transfer learning, PEFT
Fig. 1: Comparison of fusion paradigms. (a) Prior works rely on a rigid, static alignment between corresponding feature blocks. (b) Our Architectural Co-Design enables a flexible fusion policy by first enriching visual features with local priors (Conv-LoRA) and then dynamically generating tailored text descriptors for each visual level (DFG).
Fig. 2: Overview of the proposed ACD-CLIP architecture. (a) The Overall Framework: Vision and text encoders are partitioned into hierarchical Groups. Each vision group is enhanced by a trainable Conv-LoRA Adapter to instill local priors. The Dynamic Fusion Gateway (DFG) then uses each visual feature $V_{i}$ to generate a tailored text descriptor for producing a level-specific anomaly map $M_{i}$. (b) The Conv-LoRA Adapter: Our parameter-efficient adapter features a multi-branch design with multi-scale convolutions inside a LoRA bottleneck.
1 Introduction
Zero-Shot Anomaly Detection (ZSAD) leverages Vision-Language Models (VLMs) [1, 2] like CLIP [3] to bypass the need for extensive, category-specific training data required by traditional methods [4, 5]. The dominant paradigm adapts these VLMs via text prompts, which has evolved from hand-crafted ensembles (e.g., WinCLIP [6]) to learnable prompts tuned on auxiliary data. These advanced methods explore object-agnostic semantics (AnomalyCLIP [7]), dynamic prompt generation (AdaCLIP [8]), and multi-layer feature queries (CLIP-AD [9]). While significantly advancing prompt engineering, this line of work faces two fundamental limitations.
First, these methods rely on an inflexible fusion paradigm [Figure 1(a)], treating the VLM’s architecture as a fixed black box. This static, layer-wise alignment assumes a fixed semantic correspondence that often fails to capture the diverse nature of anomalies. Second, a deeper representational adaptation gap exists, as CLIP’s Vision Transformer (ViT) [10] architecture lacks the local inductive biases inherent in CNNs [11, 12], which are critical for fine-grained spatial tasks.
To address these core issues, we introduce the Architectural Co-Design CLIP (ACD-CLIP) framework [Figure 1(b)], a novel approach that synergistically refines both feature representation and cross-modal fusion. The synergy is crucial: our adapter provides the fine-grained visual details that static fusion would overlook, while the gateway supplies the adaptive mechanism required to fully exploit them. Our architectural modifications are enabled by Parameter-Efficient Fine-Tuning (PEFT) [13, 14]. While recent works have shown the value of integrating convolutional structures into PEFT adapters for general vision tasks [15, 16], our work is the first to co-design such an adapter with a dynamic fusion mechanism specifically for ZSAD. Our contributions are:
  • A parameter-efficient Conv-LoRA Adapter to inject local inductive biases for fine-grained representation.
  • A Dynamic Fusion Gateway, a vision-guided mechanism enabling a flexible, bidirectional fusion policy.
  • Validation of our synergistic co-design, which yields significant performance gains on diverse ZSAD benchmarks.
2 Methodology
As illustrated in Figure 2, our ACD-CLIP framework introduces two core innovations to adapt VLMs for ZSAD: (1) an architectural co-design for decoupled representation learning, and (2) a dynamic gateway for cross-modal fusion. We detail these components below.
2.1 Hierarchical Feature Adaptation with Local Priors
To mitigate feature entanglement, we partition the vision and text encoders into $N$ sequential Groups, encouraging each to specialize in representations at a distinct semantic level. To instill local inductive priors, we integrate a parameter-efficient Conv-LoRA Adapter into each vision group. The adapter processes the sequential output from a ViT block, $X_{\text{in}}\in\mathbb{R}^{B\times L\times C}$.
The adapter features a parallel multi-branch design ($3\times 3$ and $5\times 5$ convolutions) operating within a LoRA bottleneck. This multi-scale convolutional design allows the adapter to capture local patterns at varied receptive fields, which is crucial for identifying anomalies of different sizes. For each branch $k$, the input $X_{\text{in}}$ is projected to a low-rank tensor $X_{\text{proj}}$, reshaped into a spatial format ($\mathcal{R}_{\text{2D}}$), and passed through a convolutional bottleneck comprising a strided convolution ($\text{Conv}_{\text{down}}^{(k)}$) and a transposed convolution ($\text{Conv}_{\text{up}}^{(k)}$).
$X_{\text{proj}}=X_{\text{in}}W_{\text{down}}$ (1)
$X_{\text{up}}^{(k)}=\frac{1}{k^{2}}\operatorname{Conv}_{\text{up}}^{(k)}(\operatorname{Conv}_{\text{down}}^{(k)}(\mathcal{R}_{\text{2D}}(X_{\text{proj}})))$ (2)
The output $X_{\text{up}}^{(k)}$ is reshaped back to sequence format ($\mathcal{R}_{\text{seq}}$) and projected by $W_{\text{up}}$. The outputs from all branches are then concatenated and fused by a $1\times 1$ convolution to generate the residual update $\Delta X$, which enhances the main vision stream.
$X_{\text{branch\_out}}^{(k)}=\mathcal{R}_{\text{seq}}(X_{\text{up}}^{(k)})W_{\text{up}}$ (3)
$X_{\text{cat}}=\operatorname{Concat}(X_{\text{branch\_out}}^{(3)},X_{\text{branch\_out}}^{(5)})$ (4)
$\Delta X=\mathcal{R}_{\text{seq}}(\operatorname{Conv}_{1\times 1}(\mathcal{R}_{\text{2D}}(X_{\text{cat}})))$ (5)
2.2 Dynamic Fusion Gateway
To overcome static fusion, our Dynamic Fusion Gateway (DFG) employs a vision-guided mechanism to generate custom semantic descriptors $(T_{i}^{N},T_{i}^{A})$ for each vision feature group $V_{i}$. A global context vector is extracted from $V_{i}$ and passed through an $\text{MLP}_{\text{gate}}$ to produce logits over all $N$ text feature levels. This context vector acts as a query to dynamically weigh the importance of all text levels, effectively creating a vision-guided attention mechanism over the text features. These logits are normalized via Softmax to generate dynamic weight vectors, $\bm{\omega}_{i}^{N}$ and $\bm{\omega}_{i}^{A}$:
$\bm{\omega}_{i}^{N}=[\omega_{i,1}^{N},\dots,\omega_{i,N}^{N}]=\text{Softmax}(L_{i}^{N})$ (6)
$\bm{\omega}_{i}^{A}=[\omega_{i,1}^{A},\dots,\omega_{i,N}^{A}]=\text{Softmax}(L_{i}^{A})$ (7)
These weights dynamically fuse the set of multi-level text features $\{(T_{j}^{N},T_{j}^{A})\}_{j=1}^{N}$ to produce context-aware ”normal” ($T_{i}^{N}$) and ”abnormal” ($T_{i}^{A}$) descriptors:
$T_{i}^{N}=\sum_{j=1}^{N}\omega_{i,j}^{N}\cdot T_{j}^{N}$ (8)
$T_{i}^{A}=\sum_{j=1}^{N}\omega_{i,j}^{A}\cdot T_{j}^{A}$ (9)
The final anomaly map is the average of level-specific maps $\{M_{i}\}$, each computed from the cosine similarity between $V_{i}$ and its corresponding descriptor pair.
2.3 Training Objective and Inference
The framework is trained end-to-end with a composite loss function.
Segmentation Loss. Level-specific anomaly maps $\{M_{i}\}$ are averaged to form the final prediction map $M_{\text{seg}}$, which is optimized using a combined Focal and Dice loss $\mathcal{L}_{\text{seg}}$ against the ground-truth mask $M_{\text{GT}}$:
$\mathcal{L}_{\text{seg}}=\lambda_{\text{F}}L_{\text{F}}(M_{\text{seg}},M_{\text{GT}})+\lambda_{\text{D}}L_{\text{D}}(M_{\text{seg}},M_{\text{GT}})$ (10)
Classification Loss. For image-level classification, we compute the cosine similarity between the final global visual feature ($V_{N}^{\text{cls}}$) and the original, unfused text features from the last group ($T_{N}^{N},T_{N}^{A}$), supervised via Cross-Entropy loss $\mathcal{L}_{\text{cls}}$. We use the original text features here to provide a stable, high-level semantic anchor for the classification task, decoupling it from the fine-grained fusion used for segmentation.
Total Objective. The final training objective is a weighted sum of the two losses:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{seg}}+\lambda_{\text{cls}}\mathcal{L}_{\text{cls}}$ (11)
where $\lambda_{\text{F}},\lambda_{\text{D}},\text{ and }\lambda_{\text{cls}}$ are balancing hyperparameters.
Inference. A single forward pass generates the anomaly map $M_{\text{seg}}$ and a classification score. The final image-level score combines the direct classification output with the maximum value from $M_{\text{seg}}$.
Table 1: Quantitative comparison with state-of-the-art methods across diverse ZSAD benchmarks. Results are reported as (AUROC, AP) in percentage. Blue: Best result. Red: Second-best result.
Domain (Metric) Dataset WinCLIP CLIP-AD AnomalyCLIP AdaCLIP ours
$N=2$ $N=3$ $N=4$ $N=6$
Industrial (Pixel-level) MVTec-AD (85.1, 18.0) (89.8, 40.0) (91.1, 34.5) (86.8, 38.1) (91.7, 44.1) (91.4, 43.6) (90.9, 44.0) (90.0, 43.1)
BTAD (71.4, 11.2) (93.1, 46.7) (93.3, 42.0) (87.7, 36.6) (96.3, 51.2) (95.9, 51.2) (96.5, 51.5) (94.6, 51.1)
MPDD (95.2, 28.1) (95.1, 28.4) (96.2, 28.9) (96.6, 29.1) (97.0, 29.8) (96.3, 30.3) (96.1, 29.4) (96.6, 30.3)
RSDD (95.1, 2.1) (99.2, 31.9) (99.1, 19.1) (99.5, 38.2) (99.1, 40.7) (99.4, 41.3) (98.9, 40.4) (98.6, 40.1)
VisA (79.6, 5.0) (95.0, 26.3) (95.4, 21.3) (95.1, 29.2) (95.7, 27.8) (95.9, 29.6) (94.6, 27.5) (94.1, 27.3)
Medical (Pixel-level) ColonDB (64.8, 14.3) (80.3, 23.7) (81.9, 31.3) (79.3, 26.2) (85.0, 35.9) (85.1, 32.6) (83.3, 35.4) (80.7, 31.1)
ClinicDB (70.7, 19.4) (85.8, 39.0) (85.9, 42.2) (84.3, 36.0) (90.4, 53.5) (89.2, 54.0) (89.8, 56.1) (85.0, 49.1)
Kvasir (69.8, 27.5) (82.5, 46.2) (81.8, 42.5) (79.4, 43.8) (88.5, 60.7) (88.8, 60.2) (88.8, 61.3) (83.3, 54.8)
BrainMRI (86.0, 49.2) (96.4, 54.2) (95.6, 53.1) (93.9, 52.3) (96.6, 55.6) (95.3, 53.0) (97.0, 61.0) (96.9, 56.1)
Liver CT (96.2, 7.2) (95.4, 7.1) (93.9, 5.7) (94.5, 5.9) (97.3, 8.8) (97.2, 7.5) (96.0, 6.8) (95.3, 6.2)
Retina OCT (80.6, 39.8) (90.9, 48.7) (92.6, 55.3) (88.5, 47.1) (96.1, 66.2) (93.7, 50.9) (91.3, 48.2) (91.5, 48.7)
Industrial (Image-level) MVTec-AD (89.3, 92.9) (89.8, 95.3) (90.3, 95.1) (90.7, 95.2) (90.9, 95.7) (90.7, 95.8) (92.4, 96.8) (90.7, 95.7)
BTAD (83.3, 84.1) (85.8, 85.2) (89.1, 91.1) (91.6, 92.4) (93.5, 96.0) (94.9, 98.0) (93.3, 94.0) (95.4, 98.2)
MPDD (63.6, 71.2) (74.5, 77.9) (73.7, 77.1) (72.1, 76.9) (78.1, 83.7) (77.6, 82.3) (74.7, 79.0) (74.8, 78.2)
RSDD (85.3, 65.3) (88.3, 73.9) (73.5, 55.0) (89.1, 70.8) (94.0, 92.9) (94.3, 92.7) (93.4, 92.2) 92.9, 91.9
VisA (78.1, 77.5) (79.8, 84.3) (82.1, 85.4) (83.0, 84.9) (85.6, 88.5) (85.5, 88.1) (83.0, 86.0) (84.1, 86.7)
Medical (Image-level) BrainMRI (82.0, 90.7) (82.8, 85.5) (86.1, 92.3) (84.9, 94.2) (89.1, 97.2) (86.8, 96.9) (88.1, 97.3) (87.3, 97.1)
Liver CT (64.2, 55.9) (62.7, 51.6) (61.6, 53.1) (64.2, 56.7) (60.2, 54.2) (65.8, 55.3) (64.4, 57.3) (68.4, 58.9)
Retina OCT (42.5, 50.9) (67.9, 71.3) (75.7, 77.4) (82.7, 80.3) (84.4, 85.6) (81.1, 80.9) (80.3, 79.1) (82.0, 79.7)
3 Experiments
We conduct extensive experiments to validate the effectiveness of our proposed Architectural Co-Design with Dynamic Fusion (ACD-CLIP). We benchmark our method against state-of-the-art (SOTA) approaches on multiple industrial and medical datasets and perform detailed ablation studies to analyze the contribution of each component.
3.1 Experimental Setup
Datasets and Baselines. We evaluate on a wide range of public benchmarks. For the industrial domain, we use MVTec-AD [17], BTAD [18], MPDD [19], RSDD [20], and VisA [21]. For the medical domain, we use BrainMRI, Liver CT, and Retina OCT from BMAD [22], ColonDB [23], ClinicDB [24], Kvasir [25] to demonstrate cross-domain generalization. We compare our method against several recent SOTA methods, including WinCLIP [6], CLIP-AD [9], AnomalyCLIP [7], and AdaCLIP [8]. To maintain a strict zero-shot setting, our model is trained on the VisA dataset for all benchmarks except VisA itself, as there is no category overlap. For evaluation on the VisA benchmark, the model is trained exclusively on the MVTec-AD dataset.
Metrics. We follow standard evaluation protocols, reporting the Area Under the Receiver Operating Characteristic curve (AUROC) and Average Precision (AP) for both Pixel-level segmentation and Image-level classification tasks.
3.2 Main Results
Our method demonstrates superior performance across diverse ZSAD benchmarks. As shown in Table 1, on the MVTec-AD benchmark, our model achieves a pixel-level AUROC/AP of 91.7%/44.0% and an image-level AUROC of 92.4%, significantly outperforming prior methods like AnomalyCLIP (91.1%/34.5% pixel; 90.3% image). This advantage is consistent across other industrial datasets such as BTAD and VisA. Crucially, our approach demonstrates strong generalization to the medical domain. For instance, on ClinicDB, our model obtains a 90.4% pixel-level AUROC, surpassing AnomalyCLIP (85.9%) by a large margin. Qualitative results (Figure 3) corroborate these findings, showing our model generates precise anomaly maps that align closely with the ground truth, exhibiting higher fidelity and less background noise than competing methods.
Fig. 3: Qualitative comparison on diverse industrial and medical datasets, showing our method’s superior localization accuracy and noise suppression over state-of-the-art baselines.
Table 2: Ablation study of our proposed components. We report the average AUROC (%) for both pixel-level and image-level anomaly detection.
Configuration Avg. AUROC
Pixel-Level Image-Level
1. Baseline 82.3 81.2
2. + Conv-LoRA Adapter 89.1 (+6.8) 84.1 (+2.9)
3. + DFG 87.6 (+5.3) 85.9 (+4.7)
4. Ours (ACD-CLIP) 91.7 90.9
3.3 Ablation and Analysis
To isolate the effects of our components, we conduct an ablation study on the MVTEC-AD dataset with a two-group configuration ($N=2$). We validate our two core contributions, the Conv-LoRA Adapter and the Dynamic Fusion Gateway (DFG), against a baseline that uses a standard linear adapter and static fusion. As shown in Table 2, both components individually yield significant improvements. Specifically, adding the Conv-LoRA Adapter boosts the pixel-level AUROC by 6.8% (to 89.1%), confirming the benefit of injecting local inductive biases for dense prediction. Concurrently, the DFG enhances the image-level AUROC by 4.7% (to 85.9%), demonstrating the effectiveness of its vision-guided, context-aware fusion. Our full model, which integrates both components, achieves the highest performance (91.7% pixel-level and 90.9% image-level AUROC), proving a strong synergistic effect between the two modules. These results validate that both components are essential and complementary for achieving a substantial performance uplift.
4 CONCLUSION
In this work, we address the core limitations of Pre-trained Vision-Language Models for Zero-Shot Anomaly Detection by introducing a novel Architectural Co-Design framework. Our approach enhances both feature representation and cross-modal fusion: by injecting local inductive biases via a parameter-efficient Conv-LoRA adapter and replacing static alignment with a vision-guided Dynamic Fusion Gateway, our model learns fine-grained, context-aware features. This synergistic design achieves highly competitive results across diverse industrial and medical benchmarks, demonstrating substantial improvements in performance and stability. Ultimately, this work validates that a co-design of representation and fusion is a critical strategy for effectively adapting foundation models to dense perception tasks. We believe this principle can be extended to other domains such as zero-shot semantic segmentation and open-vocabulary detection.

### v2

@@
 2: Overview of the proposed ACD-CLIP architecture.
-(a) The Overall Framework: Vision and text encoders are partitioned into hierarchical Groups.
+(a) The Overall Framework: We structure CLIP’s vision and text encoders into a hierarchy NN sequential Groups (as illustrated, $N=3$).
 Each vision group is enhanced by a trainable Conv-LoRA Adapter to instill local priors.
@@
 While recent works have shown the value of integrating convolutional structures into PEFT adapters for general vision tasks [15, 16], our work is the first to co-design such an adapter with a dynamic fusion mechanism specifically for ZSAD.
-Our contributions are:
-  • A parameter-efficient Conv-LoRA Adapter to inject local inductive biases for fine-grained representation.
-  • A Dynamic Fusion Gateway, a vision-guided mechanism enabling a flexible, bidirectional fusion policy.
-  • Validation of our synergistic co-design, which yields significant performance gains on diverse ZSAD benchmarks.
+Our contributions are: 1) A parameter-efficient Conv-LoRA Adapter to inject local inductive biases for fine-grained representation.
+2) A Dynamic Fusion Gateway, a vision-guided mechanism enabling a flexible, bidirectional fusion policy.
+3) Validation of our synergistic co-design, which yields significant performance gains on diverse ZSAD benchmarks.
 2 Methodology
@@
 The adapter processes the sequential output from a ViT block, $X_{\text{in}}\in\mathbb{R}^{B\times L\times C}$.
-The adapter features a parallel multi-branch design ($3\times 3$ and $5\times 5$ convolutions) operating within a LoRA bottleneck.
-This multi-scale convolutional design allows the adapter to capture local patterns at varied receptive fields, which is crucial for identifying anomalies of different sizes.
-For each branch $k$, the input $X_{\text{in}}$ is projected to a low-rank tensor $X_{\text{proj}}$, reshaped into a spatial format ($\mathcal{R}_{\text{2D}}$), and passed through a convolutional bottleneck comprising a strided convolution ($\text{Conv}_{\text{down}}^{(k)}$) and a transposed convolution ($\text{Conv}_{\text{up}}^{(k)}$).
-$X_{\text{proj}}=X_{\text{in}}W_{\text{down}}$ (1)
-$X_{\text{up}}^{(k)}=\frac{1}{k^{2}}\operatorname{Conv}_{\text{up}}^{(k)}(\operatorname{Conv}_{\text{down}}^{(k)}(\mathcal{R}_{\text{2D}}(X_{\text{proj}})))$ (2)
-The output $X_{\text{up}}^{(k)}$ is reshaped back to sequence format ($\mathcal{R}_{\text{seq}}$) and projected by $W_{\text{up}}$.
-The outputs from all branches are then concatenated and fused by a $1\times 1$ convolution to generate the residual update $\Delta X$, which enhances the main vision stream.
-$X_{\text{branch\_out}}^{(k)}=\mathcal{R}_{\text{seq}}(X_{\text{up}}^{(k)})W_{\text{up}}$ (3)
-$X_{\text{cat}}=\operatorname{Concat}(X_{\text{branch\_out}}^{(3)},X_{\text{branch\_out}}^{(5)})$ (4)
-$\Delta X=\mathcal{R}_{\text{seq}}(\operatorname{Conv}_{1\times 1}(\mathcal{R}_{\text{2D}}(X_{\text{cat}})))$ (5)
+The adapter features a parallel multi-branch design, where each branch employs a distinct convolutional kernel size $k\times k$ for $k\in\{3,5\}$, operating within a LoRA bottleneck.
+This allows it to capture local patterns at varied receptive fields.
+The entire operation of the adapter, denoted as $\Phi_{\text{Adapter}}$, is formulated to generate a residual update $\Delta X$ as follows:
+$\Delta X$ $=\Phi_{\text{Adapter}}(X_{\text{in}})$ (1)
+$=\mathcal{R}_{\text{seq}}\Big{(}\operatorname{Conv}_{1\times 1}\big{(}\mathcal{R}_{\text{2D}}\big{(}\operatorname{Concat}\big{[}X_{\text{branch\_out}}^{(k)}\big{]}\big{)}\big{)}\Big{)}$
+The output of each branch, $X_{\text{branch\_out}}^{(k)}$, is generated in a two-step process.
+First, the input tensor $X_{\text{in}}$ is passed through the LoRA down-projection and a strided down-convolution to produce an intermediate downsampled feature map, $X_{\text{down}}^{(k)}$:
+$X_{\text{down}}^{(k)}=\operatorname{Conv}_{\text{down}}^{(k)}\left(\mathcal{R}_{\text{2D}}(X_{\text{in}}W_{\text{down}})\right)$ (2)
+This feature map is then upsampled via a transposed convolution, reshaped, and passed through the LoRA up-projection matrix to yield the final branch output:
+$X_{\text{branch\_out}}^{(k)}=\mathcal{R}_{\text{seq}}\left(\frac{1}{k^{2}}\operatorname{Conv}_{\text{up}}^{(k)}\left(X_{\text{down}}^{(k)}\right)\right)W_{\text{up}}$ (3)
+Here, $W_{\text{down}}$ and $W_{\text{up}}$ are the LoRA projection matrices.
+$\mathcal{R}_{\text{2D}}$ and $\mathcal{R}_{\text{seq}}$ are operators for reshaping between sequence and spatial formats.
+$\operatorname{Conv}_{\text{down}}^{(k)}$ is a strided convolution and $\operatorname{Conv}_{\text{up}}^{(k)}$ is a transposed convolution.
 2.2 Dynamic Fusion Gateway
-To overcome static fusion, our Dynamic Fusion Gateway (DFG) employs a vision-guided mechanism to generate custom semantic descriptors $(T_{i}^{N},T_{i}^{A})$ for each vision feature group $V_{i}$.
-A global context vector is extracted from $V_{i}$ and passed through an $\text{MLP}_{\text{gate}}$ to produce logits over all $N$ text feature levels.
-This context vector acts as a query to dynamically weigh the importance of all text levels, effectively creating a vision-guided attention mechanism over the text features.
-These logits are normalized via Softmax to generate dynamic weight vectors, $\bm{\omega}_{i}^{N}$ and $\bm{\omega}_{i}^{A}$:
-$\bm{\omega}_{i}^{N}=[\omega_{i,1}^{N},\dots,\omega_{i,N}^{N}]=\text{Softmax}(L_{i}^{N})$ (6)
-$\bm{\omega}_{i}^{A}=[\omega_{i,1}^{A},\dots,\omega_{i,N}^{A}]=\text{Softmax}(L_{i}^{A})$ (7)
-These weights dynamically fuse the set of multi-level text features $\{(T_{j}^{N},T_{j}^{A})\}_{j=1}^{N}$ to produce context-aware ”normal” ($T_{i}^{N}$) and ”abnormal” ($T_{i}^{A}$) descriptors:
-$T_{i}^{N}=\sum_{j=1}^{N}\omega_{i,j}^{N}\cdot T_{j}^{N}$ (8)
-$T_{i}^{A}=\sum_{j=1}^{N}\omega_{i,j}^{A}\cdot T_{j}^{A}$ (9)
-The final anomaly map is the average of level-specific maps $\{M_{i}\}$, each computed from the cosine similarity between $V_{i}$ and its corresponding descriptor pair.
+To overcome static fusion, our Dynamic Fusion Gateway (DFG) generates custom semantic descriptors for each visual feature group $V_{i}$.
+A global context vector $v_{i}^{\text{global}}=\operatorname{GAP}(V_{i})$ is extracted and passed through a gating MLP to produce logits $\bm{l}_{i}^{s}$ for each semantic state $s\in\{N,A\}$ (Normal, Abnormal).
+These logits determine the dynamic fusion weights $\omega_{i,j}^{s}$ for the set of all text features $\{T_{j}^{s}\}_{j=1}^{N}$:
+$\omega_{i,j}^{s}=\frac{\exp(l_{i,j}^{s})}{\sum_{k=1}^{N}\exp(l_{i,k}^{s})},\quad\text{where }\bm{l}_{i}^{s}=\text{MLP}_{\text{gate}}^{s}(v_{i}^{\text{global}})$ (4)
+These weights dynamically fuse the multi-level text features to produce context-aware normal ($T_{i}^{N}$) and abnormal ($T_{i}^{A}$) descriptors:
+$T_{i}^{s}=\sum_{j=1}^{N}\omega_{i,j}^{s}\cdot T_{j}^{s},\quad\text{for }s\in\{N,A\}$ (5)
+The resulting level-specific anomaly map, $M_{i}$, is then computed by applying a softmax function over the patch-wise cosine similarities between the visual features $V_{i}$ and their corresponding dynamic text descriptors:
+$M_{i}=\frac{\exp(\text{sim}(V_{i},T_{i}^{A})/\tau)}{\exp(\text{sim}(V_{i},T_{i}^{N})/\tau)+\exp(\text{sim}(V_{i},T_{i}^{A})/\tau)}$ (6)
+where $\text{sim}(\cdot,\cdot)$ denotes the cosine similarity operator and $\tau$ is a temperature parameter. The final anomaly map is the average of these level-specific maps $\{M_{i}\}$.
 2.3 Training Objective and Inference
-The framework is trained end-to-end with a composite loss function.
+The framework is trained end-to-end with a composite loss function. which jointly addresses pixel-level segmentation and image-level classification.
 Segmentation Loss.
-Level-specific anomaly maps $\{M_{i}\}$ are averaged to form the final prediction map $M_{\text{seg}}$, which is optimized using a combined Focal and Dice loss $\mathcal{L}_{\text{seg}}$ against the ground-truth mask $M_{\text{GT}}$:
-$\mathcal{L}_{\text{seg}}=\lambda_{\text{F}}L_{\text{F}}(M_{\text{seg}},M_{\text{GT}})+\lambda_{\text{D}}L_{\text{D}}(M_{\text{seg}},M_{\text{GT}})$ (10)
+The final prediction map is the average of level-specific maps, $M_{\text{seg}}=\frac{1}{N}\sum_{i=1}^{N}M_{i}$.
+It is optimized against the ground-truth mask $M_{\text{GT}}$ using a combined Focal [17] and Dice loss [18]:
+$\mathcal{L}_{\text{seg}}=\lambda_{\text{Focal}}\mathcal{L}_{\text{Focal}}(M_{\text{seg}},M_{\text{GT}})$
+$+\lambda_{\text{Dice}}\mathcal{L}_{\text{Dice}}(M_{\text{seg}},M_{\text{GT}})$ (7)
+where $\lambda_{\text{Focal}}$ and $\lambda_{\text{Dice}}$ are balancing hyperparameters.
 Classification Loss.
-For image-level classification, we compute the cosine similarity between the final global visual feature ($V_{N}^{\text{cls}}$) and the original, unfused text features from the last group ($T_{N}^{N},T_{N}^{A}$), supervised via Cross-Entropy loss $\mathcal{L}_{\text{cls}}$.
-We use the original text features here to provide a stable, high-level semantic anchor for the classification task, decoupling it from the fine-grained fusion used for segmentation.
-Total Objective. The final training objective is a weighted sum of the two losses:
+For image-level classification, we compute the cosine similarity between the final global visual feature and the original, unfused text features from the last group, supervised via Cross-Entropy loss $\mathcal{L}_{\text{cls}}$.
+We use the original text features here to provide a stable, high-level semantic anchor for the classification task.
+Total Objective. The final training objective is a weighted sum of the two losses, balanced by the hyperparameter $\lambda_{\text{cls}}$:
 $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{seg}}+\lambda_{\text{cls}}\mathcal{L}_{\text{cls}}$ (8)
-where $\lambda_{\text{F}},\lambda_{\text{D}},\text{ and }\lambda_{\text{cls}}$ are balancing hyperparameters.
 Inference.
@@
 MPDD (63.6, 71.2) (74.5, 77.9) (73.7, 77.1) (72.1, 76.9) (78.1, 83.7) (77.6, 82.3) (74.7, 79.0) (74.8, 78.2)
-RSDD (85.3, 65.3) (88.3, 73.9) (73.5, 55.0) (89.1, 70.8) (94.0, 92.9) (94.3, 92.7) (93.4, 92.2) 92.9, 91.9
+RSDD (85.3, 65.3) (88.3, 73.9) (73.5, 55.0) (89.1, 70.8) (94.0, 92.9) (94.3, 92.7) (93.4, 92.2) (92.9, 91.9)
 VisA (78.1, 77.5) (79.8, 84.3) (82.1, 85.4) (83.0, 84.9) (85.6, 88.5) (85.5, 88.1) (83.0, 86.0) (84.1, 86.7)
@@
 3 Experiments
-We conduct extensive experiments to validate the effectiveness of our proposed Architectural Co-Design with Dynamic Fusion (ACD-CLIP).
-We benchmark our method against state-of-the-art (SOTA) approaches on multiple industrial and medical datasets and perform detailed ablation studies to analyze the contribution of each component.
+To demonstrate the efficacy and robustness of our ACD-CLIP framework, we conducted a comprehensive evaluation.
+This involved benchmarking our model against state-of-the-art (SOTA) competitors across a wide range of datasets and performing in-depth ablation studies to quantify the contribution of each individual component.
 3.1 Experimental Setup
 Datasets and Baselines.
-We evaluate on a wide range of public benchmarks.
-For the industrial domain, we use MVTec-AD [17], BTAD [18], MPDD [19], RSDD [20], and VisA [21].
-For the medical domain, we use BrainMRI, Liver CT, and Retina OCT from BMAD [22], ColonDB [23], ClinicDB [24], Kvasir [25] to demonstrate cross-domain generalization.
-We compare our method against several recent SOTA methods, including WinCLIP [6], CLIP-AD [9], AnomalyCLIP [7], and AdaCLIP [8].
-To maintain a strict zero-shot setting, our model is trained on the VisA dataset for all benchmarks except VisA itself, as there is no category overlap.
-For evaluation on the VisA benchmark, the model is trained exclusively on the MVTec-AD dataset.
-Metrics.
-We follow standard evaluation protocols, reporting the Area Under the Receiver Operating Characteristic curve (AUROC) and Average Precision (AP) for both Pixel-level segmentation and Image-level classification tasks.
-3.2 Main Results
-Our method demonstrates superior performance across diverse ZSAD benchmarks.
-As shown in Table 1, on the MVTec-AD benchmark, our model achieves a pixel-level AUROC/AP of 91.7%/44.0% and an image-level AUROC of 92.4%, significantly outperforming prior methods like AnomalyCLIP (91.1%/34.5% pixel; 90.3% image).
-This advantage is consistent across other industrial datasets such as BTAD and VisA.
-Crucially, our approach demonstrates strong generalization to the medical domain.
-For instance, on ClinicDB, our model obtains a 90.4% pixel-level AUROC, surpassing AnomalyCLIP (85.9%) by a large margin.
-Qualitative results (Figure 3) corroborate these findings, showing our model generates precise anomaly maps that align closely with the ground truth, exhibiting higher fidelity and less background noise than competing methods.
+We perform a comprehensive evaluation on 12 public benchmarks spanning the industrial domain (MVTec-AD [19], VisA [20], BTAD [21], MPDD [22], RSDD [23]) and the medical domain (BrainMRI, Liver CT, and Retina OCT from BMAD [24], ColonDB [25], ClinicDB [26], Kvasir [27]).
+We benchmark against recent SOTA methods, including WinCLIP [6], CLIP-AD [9], AnomalyCLIP [7], and AdaCLIP [8].
+Implementation Details.
+Our framework is built upon the pre-trained CLIP model with a ViT-L/14 backbone at a 336px resolution.
+To maintain a strict zero-shot setting, we train our model on an auxiliary dataset with no category overlap (VisA for all non-VisA benchmarks, and MVTec-AD for the VisA benchmark).
+We report the standard Area Under the Receiver Operating Characteristic curve (AUROC) and Average Precision (AP) for both pixel-level and image-level tasks.
+3.2 Main Results and Analysis
+As shown in Table 1, our ACD-CLIP framework consistently establishes a new state-of-the-art across both industrial and medical domains.
+On the widely-used MVTec-AD benchmark, our model demonstrates a clear advantage in fine-grained localization, outperforming methods like AnomalyCLIP in pixel-level AP by a significant margin of nearly +10%.
+This superiority in capturing precise anomaly boundaries is a direct result of our architectural co-design.
+Crucially, our approach shows strong cross-domain generalization.
+When trained solely on industrial data, ACD-CLIP excels on challenging medical benchmarks, achieving a pixel-level AUROC of 90.4% on ClinicDB and 96.6% on BrainMRI.
+These quantitative gains are supported by qualitative results (Figure 3), which show our model generates significantly cleaner and more accurate anomaly maps with less background noise compared to other baselines.
 Fig. 3: Qualitative comparison on diverse industrial and medical datasets, showing our method’s superior localization accuracy and noise suppression over state-of-the-art baselines.
-Table 2: Ablation study of our proposed components. We report the average AUROC (%) for both pixel-level and image-level anomaly detection.
+Table 2: Ablation study of our proposed components on the MVTec-AD dataset with a two-group configuration ($N=2$).
+We report the average AUROC (%) for both pixel-level and image-level anomaly detection.
 Configuration Avg. AUROC
@@
 4. Ours (ACD-CLIP) 91.7 90.9
-3.3 Ablation and Analysis
-To isolate the effects of our components, we conduct an ablation study on the MVTEC-AD dataset with a two-group configuration ($N=2$).
-We validate our two core contributions, the Conv-LoRA Adapter and the Dynamic Fusion Gateway (DFG), against a baseline that uses a standard linear adapter and static fusion.
-As shown in Table 2, both components individually yield significant improvements.
-Specifically, adding the Conv-LoRA Adapter boosts the pixel-level AUROC by 6.8% (to 89.1%), confirming the benefit of injecting local inductive biases for dense prediction.
-Concurrently, the DFG enhances the image-level AUROC by 4.7% (to 85.9%), demonstrating the effectiveness of its vision-guided, context-aware fusion.
-Our full model, which integrates both components, achieves the highest performance (91.7% pixel-level and 90.9% image-level AUROC), proving a strong synergistic effect between the two modules.
-These results validate that both components are essential and complementary for achieving a substantial performance uplift.
+3.3 Ablation Study
+Our ablation study (Table 2) validates the individual and synergistic contributions of our two core components against a baseline using a linear adapter and static fusion.
+Integrating the Conv-LoRA Adapter alone boosts the pixel-level AUROC by 6.8%, confirming its critical role in providing the local inductive biases necessary for dense prediction.
+Separately, adding the Dynamic Fusion Gateway (DFG) improves the image-level AUROC by 4.7%, demonstrating the effectiveness of its vision-guided, adaptive fusion mechanism.
+The full ACD-CLIP model, which combines both, achieves the highest performance, confirming a powerful synergistic effect between enhancing feature representation and enabling dynamic cross-modal fusion.
 4 CONCLUSION

### v3

@@
 We address these limitations through an Architectural Co-Design framework that jointly refines feature representation and cross-modal fusion.
-Our method integrates a parameter-efficient Convolutional Low-Rank Adaptation (Conv-LoRA) adapter to inject local inductive biases for fine-grained representation, and introduces a Dynamic Fusion Gateway (DFG) that leverages visual context to adaptively modulate text prompts, enabling a powerful bidirectional fusion.
+Our method proposes a parameter-efficient Convolutional Low-Rank Adaptation (Conv-LoRA) adapter to inject local inductive biases for fine-grained representation, and introduces a Dynamic Fusion Gateway (DFG) that leverages visual context to adaptively modulate text prompts, enabling a powerful bidirectional fusion.
 Extensive experiments on diverse industrial and medical benchmarks demonstrate superior accuracy and robustness, validating that this synergistic co-design is critical for robustly adapting foundation models to dense perception tasks.
@@
 (a) Prior works rely on a rigid, static alignment between corresponding feature blocks.
-(b) Our Architectural Co-Design enables a flexible fusion policy by first enriching visual features with local priors (Conv-LoRA) and then dynamically generating tailored text descriptors for each visual level (DFG).
+(b) Our Architectural Co-Design enables a flexible fusion policy by enriching visual features with local priors (Conv-LoRA) and then dynamically generating tailored text descriptors for each visual level (DFG).
 Fig.
 2: Overview of the proposed ACD-CLIP architecture.
-(a) The Overall Framework: We structure CLIP’s vision and text encoders into a hierarchy NN sequential Groups (as illustrated, $N=3$).
-Each vision group is enhanced by a trainable Conv-LoRA Adapter to instill local priors.
-The Dynamic Fusion Gateway (DFG) then uses each visual feature $V_{i}$ to generate a tailored text descriptor for producing a level-specific anomaly map $M_{i}$.
+(a) The Overall Framework: We structure CLIP’s vision and text encoders into a hierarchy of NN sequential Groups (as illustrated, $N=3$).
+Each vision group is enhanced by a trainable Conv-LoRA Adapter to instill local priors, while each corresponding text group incorporates a standard LoRA adapter.
+The Dynamic Fusion Gateway (DFG) then uses each visual feature $V_{i}$ to generate a tailored text descriptor for producing a level-specific anomaly map.
 (b) The Conv-LoRA Adapter: Our parameter-efficient adapter features a multi-branch design with multi-scale convolutions inside a LoRA bottleneck.
@@
 These advanced methods explore object-agnostic semantics (AnomalyCLIP [7]), dynamic prompt generation (AdaCLIP [8]), and multi-layer feature queries (CLIP-AD [9]).
-While significantly advancing prompt engineering, this line of work faces two fundamental limitations.
+However, this line of work faces two fundamental limitations.
 First, these methods rely on an inflexible fusion paradigm [Figure 1(a)], treating the VLM’s architecture as a fixed black box.
@@
 To address these core issues, we introduce the Architectural Co-Design CLIP (ACD-CLIP) framework [Figure 1(b)], a novel approach that synergistically refines both feature representation and cross-modal fusion.
-The synergy is crucial: our adapter provides the fine-grained visual details that static fusion would overlook, while the gateway supplies the adaptive mechanism required to fully exploit them.
+The synergy is crucial: the Conv-LoRA adapter provides fine-grained details that are fully exploited by the DFG’s adaptive fusion mechanism.
 Our architectural modifications are enabled by Parameter-Efficient Fine-Tuning (PEFT) [13, 14].
@@
 2) A Dynamic Fusion Gateway, a vision-guided mechanism enabling a flexible, bidirectional fusion policy.
-3) Validation of our synergistic co-design, which yields significant performance gains on diverse ZSAD benchmarks.
+3) Validation of our method, which yields significant performance gains on diverse industrial and medical ZSAD benchmarks.
 2 Methodology
 As illustrated in Figure 2, our ACD-CLIP framework introduces two core innovations to adapt VLMs for ZSAD: (1) an architectural co-design for decoupled representation learning, and (2) a dynamic gateway for cross-modal fusion.
-We detail these components below.
 2.1 Hierarchical Feature Adaptation with Local Priors
@@
 The adapter processes the sequential output from a ViT block, $X_{\text{in}}\in\mathbb{R}^{B\times L\times C}$.
-The adapter features a parallel multi-branch design, where each branch employs a distinct convolutional kernel size $k\times k$ for $k\in\{3,5\}$, operating within a LoRA bottleneck.
-This allows it to capture local patterns at varied receptive fields.
-The entire operation of the adapter, denoted as $\Phi_{\text{Adapter}}$, is formulated to generate a residual update $\Delta X$ as follows:
+The adapter features a parallel multi-branch design with distinct $k\times k$ convolutional kernels (where $k\in\{3,5\}$) that operate within a LoRA-style bottleneck.
+This structure allows it to capture local patterns at varied receptive fields while maintaining parameter efficiency.
+The entire operation, $\Phi_{\text{Adapter}}$, generates a residual update $\Delta X$ by fusing the outputs of the parallel branches:
 $\Delta X$ $=\Phi_{\text{Adapter}}(X_{\text{in}})$ (1)
@@
 The output of each branch, $X_{\text{branch\_out}}^{(k)}$, is generated in a two-step process.
-First, the input tensor $X_{\text{in}}$ is passed through the LoRA down-projection and a strided down-convolution to produce an intermediate downsampled feature map, $X_{\text{down}}^{(k)}$:
-$X_{\text{down}}^{(k)}=\operatorname{Conv}_{\text{down}}^{(k)}\left(\mathcal{R}_{\text{2D}}(X_{\text{in}}W_{\text{down}})\right)$ (2)
-This feature map is then upsampled via a transposed convolution, reshaped, and passed through the LoRA up-projection matrix to yield the final branch output:
-$X_{\text{branch\_out}}^{(k)}=\mathcal{R}_{\text{seq}}\left(\frac{1}{k^{2}}\operatorname{Conv}_{\text{up}}^{(k)}\left(X_{\text{down}}^{(k)}\right)\right)W_{\text{up}}$ (3)
-Here, $W_{\text{down}}$ and $W_{\text{up}}$ are the LoRA projection matrices.
+First, the input tensor $X_{\text{in}}$ is projected to a low-rank space by $W_{\text{down}}$, reshaped, and passed through a $k\times k$ convolution to instill local context and transform the features, producing an intermediate feature map $X_{\text{conv}}^{(k)}$:
+$X_{\text{conv}}^{(k)}=\frac{1}{k}\operatorname{Conv}_{\text{down}}^{(k)}\left(\mathcal{R}_{\text{2D}}(X_{\text{in}}W_{\text{down}})\right)$ (2)
+This feature map is then processed by a second $k\times k$ convolution for further feature refinement, reshaped back into a sequence, and projected up by $W_{\text{up}}$ to yield the final branch output:
+$X_{\text{branch\_out}}^{(k)}=\mathcal{R}_{\text{seq}}\left(\frac{1}{k}\operatorname{Conv}_{\text{up}}^{(k)}\left(X_{\text{conv}}^{(k)}\right)\right)W_{\text{up}}$ (3)
+Here, $W_{\text{down}}$ and $W_{\text{up}}$ are the LoRA projection matrices performing channel-wise compression and expansion.
 $\mathcal{R}_{\text{2D}}$ and $\mathcal{R}_{\text{seq}}$ are operators for reshaping between sequence and spatial formats.
-$\operatorname{Conv}_{\text{down}}^{(k)}$ is a strided convolution and $\operatorname{Conv}_{\text{up}}^{(k)}$ is a transposed convolution.
+$\operatorname{Conv}_{\text{up}}^{(k)}$ and $\operatorname{Conv}_{\text{down}}^{(k)}$ are standard $k\times k$ convolutions that operate on feature channels for transformation and local context infusion while preserving spatial dimensions.
 2.2 Dynamic Fusion Gateway
@@
 As shown in Table 1, our ACD-CLIP framework consistently establishes a new state-of-the-art across both industrial and medical domains.
-On the widely-used MVTec-AD benchmark, our model demonstrates a clear advantage in fine-grained localization, outperforming methods like AnomalyCLIP in pixel-level AP by a significant margin of nearly +10%.
+On the widely-used MVTec-AD benchmark, our model excels at fine-grained localization, boosting pixel-level AP by nearly +10% over AnomalyCLIP.
 This superiority in capturing precise anomaly boundaries is a direct result of our architectural co-design.
@@
 When trained solely on industrial data, ACD-CLIP excels on challenging medical benchmarks, achieving a pixel-level AUROC of 90.4% on ClinicDB and 96.6% on BrainMRI.
-These quantitative gains are supported by qualitative results (Figure 3), which show our model generates significantly cleaner and more accurate anomaly maps with less background noise compared to other baselines.
+Qualitative results (Figure 3) corroborate these gains, showing our model produces cleaner, more accurate anomaly maps with less noise.
+Interestingly, our results across different numbers of groups ($N$) show that performance generally peaks around $N=3$, suggesting a trade-off between hierarchical specialization and model complexity, where deeper partitioning does not necessarily yield further gains and may lead to overfitting on the auxiliary dataset.
 Fig. 3: Qualitative comparison on diverse industrial and medical datasets, showing our method’s superior localization accuracy and noise suppression over state-of-the-art baselines.
@@
 3.3 Ablation Study
-Our ablation study (Table 2) validates the individual and synergistic contributions of our two core components against a baseline using a linear adapter and static fusion.
+Our ablation study (Table 2) validates the individual and synergistic contributions of our two core components against a baseline using a standard LoRA adapter on ViT and static fusion.
 Integrating the Conv-LoRA Adapter alone boosts the pixel-level AUROC by 6.8%, confirming its critical role in providing the local inductive biases necessary for dense prediction.

### v4

@@
-Architectural Co-Design for Zero-Shot Anomaly Detection: Decoupling Representation and Dynamically Fusing Features in CLIP
+ACD-CLIP: Decoupling Representation and Dynamic Fusion for Zero-Shot Anomaly Detection
 Abstract
-Pre-trained Vision-Language Models (VLMs) face a significant adaptation gap when applied to Zero-Shot Anomaly Detection (ZSAD), stemming from their lack of local inductive biases for dense prediction and their reliance on inflexible feature fusion paradigms.
+Pre-trained Vision-Language Models (VLMs) struggle with Zero-Shot Anomaly Detection (ZSAD) due to a critical adaptation gap: they lack the local inductive biases required for dense prediction and employ inflexible feature fusion paradigms.
 We address these limitations through an Architectural Co-Design framework that jointly refines feature representation and cross-modal fusion.
@@
 Second, a deeper representational adaptation gap exists, as CLIP’s Vision Transformer (ViT) [10] architecture lacks the local inductive biases inherent in CNNs [11, 12], which are critical for fine-grained spatial tasks.
-To address these core issues, we introduce the Architectural Co-Design CLIP (ACD-CLIP) framework [Figure 1(b)], a novel approach that synergistically refines both feature representation and cross-modal fusion.
+To address these core issues, we introduce the Architectural Co-Design (ACD-CLIP) framework [Figure 1(b)], a novel approach that synergistically refines both feature representation and cross-modal fusion.
 The synergy is crucial: the Conv-LoRA adapter provides fine-grained details that are fully exploited by the DFG’s adaptive fusion mechanism.
@@
 To instill local inductive priors, we integrate a parameter-efficient Conv-LoRA Adapter into each vision group.
-The adapter processes the sequential output from a ViT block, $X_{\text{in}}\in\mathbb{R}^{B\times L\times C}$.
+Distinct from prior work that utilizes a Mixture-of-Experts approach for dynamic feature scaling [16], our adapter employs a parallel multi-branch architecture with distinct kernel sizes, specifically tailored to capture the varied local patterns essential for fine-grained anomaly detection.
+The adapter processes the sequential output from a ViT block, $X_{\text{in}}\in\mathbb{R}^{B\times L\times C}$, where $B$ is the batch size, $L$ is the sequence length, and $C$ is the channel dimension.
 The adapter features a parallel multi-branch design with distinct $k\times k$ convolutional kernels (where $k\in\{3,5\}$) that operate within a LoRA-style bottleneck.
@@
 $=\mathcal{R}_{\text{seq}}\Big{(}\operatorname{Conv}_{1\times 1}\big{(}\mathcal{R}_{\text{2D}}\big{(}\operatorname{Concat}\big{[}X_{\text{branch\_out}}^{(k)}\big{]}\big{)}\big{)}\Big{)}$
+where $\operatorname{Concat}[\cdot]$ aggregates the feature maps from each branch along the channel dimension, and a subsequent $1\times 1$ convolution, $\operatorname{Conv}_{1\times 1}$, adaptively fuses these multi-scale features into a unified residual update.
 The output of each branch, $X_{\text{branch\_out}}^{(k)}$, is generated in a two-step process.
@@
 2.3 Training Objective and Inference
-The framework is trained end-to-end with a composite loss function. which jointly addresses pixel-level segmentation and image-level classification.
+The framework is trained end-to-end with a composite loss function, which jointly addresses pixel-level segmentation and image-level classification.
 Segmentation Loss.
@@
 A single forward pass generates the anomaly map $M_{\text{seg}}$ and a classification score.
-The final image-level score combines the direct classification output with the maximum value from $M_{\text{seg}}$.
+The final image-level anomaly score combines the direct classification output with the maximum value from $M_{\text{seg}}$.
 Table 1: Quantitative comparison with state-of-the-art methods across diverse ZSAD benchmarks. Results are reported as (AUROC, AP) in percentage. Blue: Best result. Red: Second-best result.
@@
 As shown in Table 1, our ACD-CLIP framework consistently establishes a new state-of-the-art across both industrial and medical domains.
-On the widely-used MVTec-AD benchmark, our model excels at fine-grained localization, boosting pixel-level AP by nearly +10% over AnomalyCLIP.
+On the widely-used MVTec-AD benchmark, our model excels at fine-grained localization, boosting the pixel-level AP by nearly 10 percentage points over AnomalyCLIP.
 This superiority in capturing precise anomaly boundaries is a direct result of our architectural co-design.
-Crucially, our approach shows strong cross-domain generalization.
-When trained solely on industrial data, ACD-CLIP excels on challenging medical benchmarks, achieving a pixel-level AUROC of 90.4% on ClinicDB and 96.6% on BrainMRI.
-Qualitative results (Figure 3) corroborate these gains, showing our model produces cleaner, more accurate anomaly maps with less noise.
+Trained only on industrial data, ACD-CLIP shows strong cross-domain generalization, achieving 90.4% and 96.6% pixel-level AUROC on ClinicDB and BrainMRI, respectively.
+Qualitative results (Fig.
+3) visually substantiate these gains, showcasing cleaner and more precise anomaly maps.
 Interestingly, our results across different numbers of groups ($N$) show that performance generally peaks around $N=3$, suggesting a trade-off between hierarchical specialization and model complexity, where deeper partitioning does not necessarily yield further gains and may lead to overfitting on the auxiliary dataset.
@@
 3.3 Ablation Study
-Our ablation study (Table 2) validates the individual and synergistic contributions of our two core components against a baseline using a standard LoRA adapter on ViT and static fusion.
+Our ablation study (Table 2) validates each component’s contribution against a baseline using standard LoRA and static fusion.
 Integrating the Conv-LoRA Adapter alone boosts the pixel-level AUROC by 6.8%, confirming its critical role in providing the local inductive biases necessary for dense prediction.
 Separately, adding the Dynamic Fusion Gateway (DFG) improves the image-level AUROC by 4.7%, demonstrating the effectiveness of its vision-guided, adaptive fusion mechanism.
-The full ACD-CLIP model, which combines both, achieves the highest performance, confirming a powerful synergistic effect between enhancing feature representation and enabling dynamic cross-modal fusion.
+The full ACD-CLIP model achieves the highest performance, confirming the powerful synergy between our representation and fusion modules.
 4 CONCLUSION
@@
 This synergistic design achieves highly competitive results across diverse industrial and medical benchmarks, demonstrating substantial improvements in performance and stability.
-Ultimately, this work validates that a co-design of representation and fusion is a critical strategy for effectively adapting foundation models to dense perception tasks.
+Ultimately, this work validates that the co-design of representation and fusion is a critical strategy for effectively adapting foundation models to dense perception tasks.
 We believe this principle can be extended to other domains such as zero-shot semantic segmentation and open-vocabulary detection.

### v5

@@
 Extensive experiments on diverse industrial and medical benchmarks demonstrate superior accuracy and robustness, validating that this synergistic co-design is critical for robustly adapting foundation models to dense perception tasks.
+The source code is available at https://github.com/cockmake/ACD-CLIP.
 Index Terms—  anomaly detection, multimodal feature fusion, vision-language model, transfer learning, PEFT
@@
 Each vision group is enhanced by a trainable Conv-LoRA Adapter to instill local priors, while each corresponding text group incorporates a standard LoRA adapter.
-The Dynamic Fusion Gateway (DFG) then uses each visual feature $V_{i}$ to generate a tailored text descriptor for producing a level-specific anomaly map.
+The Dynamic Fusion Gateway then uses each visual feature $V_{i}$ to generate a tailored text descriptor for producing a level-specific anomaly map.
 (b) The Conv-LoRA Adapter: Our parameter-efficient adapter features a multi-branch design with multi-scale convolutions inside a LoRA bottleneck.
 1 Introduction
-Zero-Shot Anomaly Detection (ZSAD) leverages Vision-Language Models (VLMs) [1, 2] like CLIP [3] to bypass the need for extensive, category-specific training data required by traditional methods [4, 5].
-The dominant paradigm adapts these VLMs via text prompts, which has evolved from hand-crafted ensembles (e.g., WinCLIP [6]) to learnable prompts tuned on auxiliary data.
-These advanced methods explore object-agnostic semantics (AnomalyCLIP [7]), dynamic prompt generation (AdaCLIP [8]), and multi-layer feature queries (CLIP-AD [9]).
+Zero-Shot Anomaly Detection (ZSAD) adapts Vision-Language Models (VLMs) [1, 2] like CLIP [3] to circumvent the extensive training data required by traditional methods [4, 5].
+The dominant paradigm uses text prompts, evolving from hand-crafted ensembles to learnable prompts tuned on auxiliary data (WinCLIP [6]).
+These advanced methods have explored object-agnostic semantics (AnomalyCLIP [7]), dynamic prompt generation (AdaCLIP [8]), and multi-layer feature queries (CLIP-AD [9]).
 However, this line of work faces two fundamental limitations.
@@
 Our contributions are: 1) A parameter-efficient Conv-LoRA Adapter to inject local inductive biases for fine-grained representation.
-2) A Dynamic Fusion Gateway, a vision-guided mechanism enabling a flexible, bidirectional fusion policy.
+2) A Dynamic Fusion Gateway (DFG), a vision-guided mechanism enabling a flexible, bidirectional fusion policy.
 3) Validation of our method, which yields significant performance gains on diverse industrial and medical ZSAD benchmarks.
 2 Methodology
-As illustrated in Figure 2, our ACD-CLIP framework introduces two core innovations to adapt VLMs for ZSAD: (1) an architectural co-design for decoupled representation learning, and (2) a dynamic gateway for cross-modal fusion.
+As illustrated in Figure 2, our ACD-CLIP framework introduces two core innovations to adapt VLMs for ZSAD: (1) a Conv-LoRA Adapter to instill local inductive biases for fine-grained representation, and (2) a Dynamic Fusion Gateway for adaptive cross-modal fusion.
 2.1 Hierarchical Feature Adaptation with Local Priors
@@
 Trained only on industrial data, ACD-CLIP shows strong cross-domain generalization, achieving 90.4% and 96.6% pixel-level AUROC on ClinicDB and BrainMRI, respectively.
-Qualitative results (Fig.
-3) visually substantiate these gains, showcasing cleaner and more precise anomaly maps.
+Qualitative results (Figure 3) substantiate these gains, showcasing cleaner and more precise anomaly maps.
 Interestingly, our results across different numbers of groups ($N$) show that performance generally peaks around $N=3$, suggesting a trade-off between hierarchical specialization and model complexity, where deeper partitioning does not necessarily yield further gains and may lead to overfitting on the auxiliary dataset.